OpenStudy (anonymous):

Algebra Help Plz 3. Order the group of quadratic functions from widest to narrowest graph. y = 2x^2; y = 10x^2; y = 8x^2 A) y = 2x^2, y = 10x^2; y = 8x^2 B) y = 8x^2; y = 10x^2; y = 2x^2 C) y = 10x^2; y = 8x^2; y = 2x^2 D) y = 2x^2, y = 8x^2, y = 10x^2

OpenStudy (anonymous):

@cookienano

OpenStudy (anonymous):

The answer would be D because once the coefficient gets larger the parabola gets narrower. If we use a graphing calculator or a website like Desmos we can see that y=2x^2 is wider than y=8x^2 {which is wider than y=10x^2}.

OpenStudy (anonymous):

4. Order the group of quadratic functions from widest to narrowest graph. y = -1/3x^2; y = 2/3x^2; y = 3/2x^2 A) y = -1/3x^2; y = 2/3x^2; y = 3/2x^2 B) y = 3/2x^2; y = 2/3x^2; y = -1/3x^2 C) y = 2/3x^2; y = -1/3x^2; y = 3/2x^2 D) y = -1/3x^2; y = 3/2x^2; y = 2/3x^2 5. If an object is dropped from a height of 200 feet, the function h(t) = -16t^2 + 200 gives the height of the object after t seconds. Approximately, when will the object hit the ground? A) 200.00 seconds B) 184.00 seconds C) 3.54 seconds D) 0.78 seconds

OpenStudy (anonymous):

@cookienano

OpenStudy (anonymous):

4. The answer would be A because as I said the bigger the coefficient the narrower the parabola. 5. For this question we have the values of a and b. a = -16 and b = 200 So to find the answer we would use -(b/2a) so lets plug in the values of a and b in... -(200/[2][-16]) --> -(200/-32) --> -(-6.25) --> 6.25 seconds Looks like that isn't an answer in your choices, did you write them correctly?